Wonga Manufacturing Inc. makes high quality silver spoons. The forecast for 2020 is to make 1000 spoons. The firm’s production process is the most efficient and produces 8 spoons per day. The firms working policy is 250 days a year, it cost the firm K10 to set up the equipment and 50ngwee for carrying inventory from one period to the next.
I. Calculate the economic order quantity [5 Marks] II. Find the annual carrying cost [3 Marks] III. Find the annual setup cost [3 Marks] IV. Find the maximum number of spoon the firm can accumulate. [5 Marks] V. Find the number of production order for 2020 [4 Marks] VI. Find the number of days it take to make the next order. [5 Marks] [Total: 25 Marks]
Answer #1

Annual demand (D) = 1000

Daily demand (d) = 1000/250 = 4 per day

Daily production rate (p) = 8 per day

Setup cost (S) = 10

Holding cost (H) = 0.50

I.

Economic quantity Q = sqrt( 2DS / H(1-d/p) ) = sqrt(2*1000*10/(0.50*(1-4/8))) = 282.82 or 283 units

II.

Annual carrying cost = (HQ/2)*(1-d/p) = (0.50*283/2)*(1-4/8) = 35.37

III.

Annual setup cost = DS/Q = 1000*10/283 = 35.33

IV.

Maximum inventory = Q*(1-d/p) = 283*(1-4/8) = 141.5

V.

Number of orders = 1000/283 = 3.53 or 4 times

VI.

Number of days till next order = 250/4 = 62.5 days